OpenStudy (anonymous):

solve log (base 3) (5x-2)= 4

OpenStudy (saifoo.khan):

\[\log_3 ( 5x -2) = 4\] \[5x -2 = 3^4\] Can u solve now?

OpenStudy (saifoo.khan):

the property i used, \[\Large \log_a b = c \to b = a^c\]

OpenStudy (anonymous):

how do you get rid of the log base 3 again..

OpenStudy (saifoo.khan):

It's a rule - i dont remember it's name though. i can google if u want. but we change it the i showed in my 2nd comment.

OpenStudy (anonymous):

yeah ill look it up.. i just wanna know how to get rid of the log. I just cant remember how.. :/

OpenStudy (saifoo.khan):

It's just like this: \[\large \log_a b = c \to b = a^c\]or, \(\log_e 2 = 2\) can be written as \(2= e^2\)

OpenStudy (anonymous):

ok